OpenStudy (anonymous):

solve by substitution . 3x-4y=8 x-2y=2

OpenStudy (anonymous):

simply take one of the equations and solve for y then plug that equation into the other to get x. Once you get your x value just plug it back into the y equation that you found earlier to get your y value. If you need additional help just ask.

OpenStudy (anonymous):

could you please guide me through that?

OpenStudy (anonymous):

alright. Let's do this: Take the x-2y=2. Move the x to the other side via subtraction to get -2y= 2-x Now you need to get y by itself so divide each side by -2. You will then get y = (2-x)/-2. Now that you have this you can take this y value and plug it back into the other equation. So

OpenStudy (anonymous):

So you will have 3x-4((2-x)/-2)=8. Now solve for x. I hope you can do this... After you find the value of x. You plug it back into your equation of y = (2-x)/-2 to get your y value.

OpenStudy (anonymous):

how do i solve for x? sorry, im rally bad n math =P

OpenStudy (anonymous):

That's fine. Now ignore what I put above here is a faster way. You have your 2 equations.... Take the x-2y=2 and instead of solving for y, solve for x instead (don't have to deal with fractions then). Add the 2y over to the right to get x = 2+2y. With this in mind, plug this into the other equation of 3x-4y=8 for x. Thus you will get 3(2+2y)-4y=8. Distribute the 3 within the () to get 6+6y-4y=8. Combine your like terms by subtracting 6y-4y=2y and moving the 6 to the 8 by subtraction. When that is all said and done, you will end up with 6y=6. Simply dividing both sides by 6 to get your y value which is 1. Now that you have you y=1 plug that back into you x equation you found earlier of x=2+2y. So you will get x=2+2(1) in otherwords 4. So x=4 and y=1.

OpenStudy (anonymous):

oh that makes so much sense now =P thank you SOO MUCH =}

OpenStudy (anonymous):

Your welcome =)